OpenStudy (anonymous):

What is the base of a triangle with height = 7.4 cm and area = 48.1 cm2? 355.94 cm 6.5 cm 13 cm 177.97 cm

OpenStudy (anonymous):

Let base b =x cm h=7.4 cm A=48.1 cm^2 A=1/2 *b*h

OpenStudy (anonymous):

48.1=1/2*x*7.4=3.7x x=48.1/3.7 calculate and get the required answer.